OpenStudy (anonymous):

Explain, in complete sentences, how you would divide the following expression: tenth root of x to the seventh power over seventh root of x to the fourth power

OpenStudy (anonymous):

okay when you are taking the nth root of something you could also write it as x^(1/n) you know what I mean? So the 10th root is the same as x^(1/10). Then you just use exponent rules do you know these? x^2^2 = x^4 and x^3/x^2 = x know what I mean?

OpenStudy (anonymous):

??

OpenStudy (campbell_st):

\[\frac{(x^{\frac{1}{10}})^7}{(x^\frac{1}{7})^4}\] use the power of a power rule to simplify the powers of x in both the numerator and denominator \[(x^a)^b = x^{a \times b}\] this gives \[\frac{x^{\frac{7}{10}}}{x^{\frac{4}{7}}}\] when dividing like pronumerals subtract the powers \[x^{\frac{7}{10} - \frac{4}{7}}\] results in an answer of \[x^{\frac{9}{70}}\]

OpenStudy (anonymous):

What do you need explained? just start with the basics if you had squarert(x) what is another way you could write that?